There are coins including nickels, dimes and quarters totaling $5.45. If there are twice as many quarters as dimes and 11 more nickels than quarters, how many of each coin are there?
Answer by richwmiller(17219)

25q+5n+10d=545
q=2d
d=q/2
q+11=n
substitute q/2 for d and q+11 for n in the first equation. ans solve for q

d = 7, n = 25, q = 14